### v1

# ACORD: An Expert-Annotated Retrieval Dataset for Legal Contract Drafting

Abstract
Information retrieval, specifically contract clause retrieval, is foundational to contract drafting because lawyers rarely draft contracts from scratch; instead, they locate and revise the most relevant precedent. We introduce the Atticus Clause Retrieval Dataset (ACORD), the first retrieval benchmark for contract drafting fully annotated by experts. ACORD focuses on complex contract clauses such as Limitation of Liability, Indemnification, Change of Control, and Most Favored Nation. It includes 114 queries and over 126,000 query-clause pairs, each ranked on a scale from 1 to 5 stars. The task is to find the most relevant precedent clauses to a query. The bi-encoder retriever paired with pointwise LLMs re-rankers shows promising results. However, substantial improvements are still needed to effectively manage the complex legal work typically undertaken by lawyers. As the first retrieval benchmark for contract drafting annotated by experts, ACORD can serve as a valuable IR benchmark for the NLP community.
1 Introduction
Figure 1: A showcase of the clause retrieval process integral to the ACORD dataset. The input is a query, “Termination for Convenience,” and the output is a ranked list of the most relevant clauses from various agreements. The clauses are accompanied by their sources, thus providing critical context for legal professionals.
Contracts are the backbone of modern business, with millions created daily. A 2023 ALM and Bloomberg Law survey found that 43% of corporate counsels spend at least half their time on tasks like drafting, editing, and negotiating contracts (Bloomberg Law (2023); Pery (2019)). Lawyers rarely write contracts from scratch; they rely on finding and adapting relevant examples. High-quality precedents are crucial for handling complex clauses that involve nuanced legal concepts and require precise language. Mistakes in wording or missing key ideas can lead to disputes, liabilities, or invalid clauses.
Research shows that Large Language Models (LLMs) can effectively analyze legal contracts and identify issues within them (Hoffman and Arbel, 2023; Martin et al., 2024). However, LLMs struggle to draft contracts independently. Table 1 compares an LLM-generated clause with two top-rated precedent clauses from ACORD. Lawyer edits, shown with strike-through and underlined text, highlight two main issues: (1) the LLM uses “boilerplate” language that creates conflicts between clauses, and (2) it introduces uncommon phrasing not found in precedents.
Prompt: draft a contract clause where liability cap is based on purchase price Query: liability cap is based on purchase price
LLM-Generated Result ACORD 5-Star Precedent Clause ACORD 4-Star Precedent Clause
1. [Section Number] Cap on Liabilitya Notwithstanding any provision to the contrary in this Agreement, Except as set forth in Section [Number] below,b Tthe total aggregate liability of [Seller’s Name] to [Buyer’s Name], whether in contract, tort (including negligence), breach of statutory duty, or otherwise, arising under or in connection with this Agreement shall not exceed the total purchase priceamount paid and payable by [Buyer’s Name] to [Seller’s Name] under this Agreement (the “Purchase Price”). c 2. [Section Number]. Exclusions from Liability Cap The limitation of liability set forth in this clause Section [Number] shall not apply to: a. Liability resulting from gross negligence, willful misconduct, or fraud by [Seller’s Name]; b. Any breach of confidentiality obligations as specified in this Agreement; c. Any liability which cannot be excluded or limited by applicable law.d 3. Section [Number] No Indirect or Consequential Damages In no event shall [Seller’s Name] be liable to [Buyer’s Name] for any indirect, incidental, special, consequential, or punitive damages, including but not limited to loss of profits or revenue, loss of use, or loss of data, even if advised of the possibility of such damages. 4. Application of Limitations The parties agree that the limitations of liability set forth in this clause are fundamental elements of the basis of the bargain between [Seller’s Name] and [Buyer’s Name], and the Purchase Price reflects these limitations.e 5. Sponsor’s Maximum Obligation; Indemnification. Racing represents to the Sponsor that the Sponsor’s aggregate obligation hereunder will not exceed the amount of the sponsorship fee set forth in Section 3 hereof (or such lesser amount as is payable by the Sponsor in the event that this Agreement is terminated pursuant to Section 14 hereof), plus, if applicable, collection costs that may be reasonably incurred by Racing in a legal proceeding to collect all or any part thereof (the “Maximum Obligation”). Racing agrees to indemnify the Sponsor and its officers, directors, agents and employees and to hold them harmless from any loss, claim, cost, damage or liability in excess of the Maximum Obligation which (i) the Sponsor shall incur as a result of this Agreement, or (ii) arises from any failure by Racing to perform any of its obligations hereunder. 12. Limitation of Liability 12.1 IN NO EVENT SHALL EITHER PARTY BE LIABLE TO THE OTHER PARTY FOR ANY INCIDENTAL, CONSEQUENTIAL, INDIRECT, SPECIAL, OR PUNITIVE DAMAGES (INCLUDING, BUT NOT LIMITED TO, LOST PROFITS, BUSINESS INTERRUPTION, LOSS OF BUSINESS INFORMATION OR OTHER PECUNIARY LOSS) REGARDLESS OF WHETHER SUCH LIABILITY IS BASED ON BREACH OF CONTRACT, TORT (INCLUDING NEGLIGENCE), STRICT LIABILITY, BREACH OF WARRANTIES, FAILURE OF ESSENTIAL PURPOSE OR OTHERWISE AND EVEN IF ADVISED OF THE POSSIBILITY OF SUCH DAMAGES. 12.2 Except with regard to a breach of confidentiality, a party’s indemnification obligations hereunder, or infringement of intellectual property rights, either party’s total liability to the other party under this Agreement shall be limited to the amounts paid or payable by the Reseller to Todos during the twelve-month period preceding the interposition of the claim.
a Attorney Note: change “1.” to a placeholder because a Cap on Liability clause would never be in Section 1.
b Attorney Note: the reference to “Notwithstanding any provision to the contrary in this Agreement” is wrong because this subsection should NOT take precedence over the “Exclusions from Liability Cap” below. Therefore, it should be replaced with “Except as set forth in Section [Number] below.”
c Attorney Note: liability cap should not be at purchase price paid in the event the buyer hasn’t paid in full or fails to pay timely. ACORD 5-star and 4-star precedent clauses fix this defect by capping the liability at the amount “paid or payable” or referring to the amount in the payment section of the contract.
d Attorney’s Note: This is not wrong; however, Exceptions to Liability Cap are often included in the same sentence as the liability cap instead of being in a stand-alone subsection.
e Attorney Note: Section 4 is not commonly seen in commercial contracts. It is unclear what the intent or purpose of this clause is.
Table 1: Comparison between a clause generated by a large language model (LLM) and two precedent clauses rated 5-star and 4-star in the ACORD dataset. Deficiencies in the LLM-generated clause, such as inconsistent phrasing and omitted legal concepts, are detailed with lawyer annotations using strike-throughs and underlined text. Attorney notes in footnotes explain why the LLM-generated clause falls short.
To enhance LLMs’ contract drafting abilities, we use Retrieval-Augmented Generation (RAG), which combines LLMs’ text generation with a retrieval system. RAG mimics how lawyers improve drafting by first searching for relevant precedents and then basing their drafts on those examples. This approach has been shown to reduce hallucinations and enhance performance (Lewis et al., 2021; Niu et al., 2024; Magesh et al., 2024). By referencing high-quality, tested precedents, LLMs can produce better contract drafts. We must evaluate how well models can retrieve the relevant contract clauses to ensure RAG can work effectively.
Clause retrieval poses several challenges. Contracts are multi-layered, with sections and subsections that can span pages, often broken into paragraphs outlining obligations (what is required) and exceptions (what is not required). Cross-references within clauses add further complexity, such as defined terms (e.g., “as defined in Section 4”) or exceptions (e.g., “except as set forth in Sections 9 and 11(a)” or “notwithstanding anything to the contrary”). These references often point to sections located pages apart, yet they are crucial for understanding a clause’s meaning. Additionally, whether one precedent clause should rank higher than the others can be subjective and depends on the user’s individual experience, the industry they are in, and their unique needs, among other things. This subjectivity is reflected in the annotators’ disagreement rate of 21%. [^1] The fraction of ratings where the annotators significantly disagreed on the score or where they disagreed on the relevance.
Datasets with expert-annotated, domain-specific benchmarks are essential for improving LLMs in clause retrieval and contract drafting. However, such datasets are rare due to the high cost of expert annotation and the confidentiality of proprietary contracts. Many existing legal contract benchmarks consist of multiple-choice Q&A formats, often derived from bar exam questions, and do not reflect the work lawyers perform, with none focusing specifically on contract drafting. To address this, we introduce ACORD, the Atticus Clause Retrieval Dataset, the first expert-annotated retrieval benchmark explicitly designed for contract drafting.
ACORD addresses the challenge of retrieving complex and heavily negotiated contract clauses, such as Limitation of Liability, Indemnification, Most Favored Nation, and Termination for Convenience. These clauses are central to contract drafting and require careful negotiation and precise language. The task is to retrieve the most relevant precedent clauses for a given query, where the output is a ranked list of top-rated clauses from a large corpus. Figure 1 illustrates this retrieval process. The dataset was developed collaboratively by seasoned lawyers, student annotators, and machine learning researchers. It comprises 114 unique queries and over 126,000 query-clause pairs, each annotated with a 1-to-5-point relevance score by experts. This rigorous annotation ensures that ACORD is a robust and reliable benchmark for evaluating retrieval models in legal contexts.
ACORD provides a new framework for assessing machine learning models’ ability to retrieve information critical for contract drafting tasks. We hope that ACORD enables researchers and ML practitioners to evaluate search systems properly on a nuanced legal search task and to make progress toward tackling real-world challenges in legal drafting. The dataset and software will be made publicly available on GitHub.
2 Related Work
Despite being one of the most common legal tasks (Pery, 2019; Bloomberg Law, 2023), contract drafting and clause retrieval have been under-studied in NLP literature due to the lack of domain-specific benchmark datasets.
2.1 Information Retrieval
Information retrieval (IR) is the process of locating relevant documents in response to a user query, as seen in web search tools such as Google and Bing. Historically, lexical-based retrieval has been foundational in IR (Hambarde and Proenca, 2023), with BM25 being one of the most popular ranking functions. BM25 scores document relevance based on term frequency and inverse document frequency while mitigating the impact of prevalent words (Robertson and Zaragoza, 2009). However, lexical-based retrieval methods face the “lexical gap”, where reliance on exact word matches often fails to capture relevant semantic connections such as synonymy and word-order-dependent relationships (Thakur et al., 2021). Recent approaches have moved toward neural IR to mitigate this, designed to capture semantic connections beyond simple lexical matches. These IR models fall into two primary categories:
  • Retriever Models: These models map queries and documents into a vector space, allowing pre-computed document representations to be indexed. For example, dense retrievers based on Transformer models (Vaswani et al., 2023) have shown robust performance in many tasks such as citation prediction, argument retrieval, summarization, duplicate detection, and question-answering (Thakur et al., 2021; Muennighoff et al., 2023). Another example is sparse embeddings that learned token-level contextualized representations using Transformers that still could be efficiently indexed using an efficient inverse index (Zhao et al., 2020; Formal et al., 2021).
  • Neural Reranking Models: These models enhance initial retrieval results (often from a BM25 or efficient retrieval model) by re-ordering the returned documents. Leveraging cross-attention mechanisms in models such as BERT has significantly boosted reranking performance (Devlin et al., 2019). ColBERT (Contextualized late interaction over BERT) further refines this approach by creating token-level embeddings for both queries and documents, using a maximum-similarity mechanism to identify relevant documents. However, this comes with higher computational costs (Devlin et al., 2019; Khattab and Zaharia, 2020). Recent research highlights that instruction-tuned large language models (LLMs) can surpass traditional supervised cross-encoders in zero-shot passage reranking tasks (Sun et al., 2023). In particular, listwise, pointwise, or pairwise reranking have received the most attention (Tang et al., 2024; Zhuang et al., 2024; Qin et al., 2024).
2.2 Large Language Models (LLMs) and Retrieval-Augmented Generation (RAG)
Despite their advanced text generation capabilities, LLMs have not been widely adopted in contract drafting due to the risk of hallucination. Existing works (Wang et al., 2024; Dahl et al., 2024) show that the generations of LLMs can be unreliable, untrustworthy, and risky in many cases. Many recent studies focus on developing and enhancing various RAG techniques to reduce hallucination and improve model performance in text generation (Niu et al., 2024; Magesh et al., 2024). Given the absence of an information retrieval-specific dataset tailored to this task, the open question is whether these methodologies can effectively retrieve precedent clauses for contract drafting purposes.
2.3 Case Law and Statute Retrieval
There are several case law retrieval datasets, including LePARD, CLERC, COLLIE, ECtHR-PCR (Li and Harada, 2022; Hou et al., 2024; Rabelo et al., 2021; T.y.s.s. et al., 2024). However, the queries in these datasets are not constructed by experts. Instead, they are synthetically created by taking blocks of text from case law with the citations redacted. The document referred to in the redacted citation serves as the corresponding relevant document for the query-document pair. This methodology cannot be used to create a contract IR dataset because contract clauses do not inherently include references or citations that can be used to infer query-document pairs. Legal experts need to generate queries and ranked clauses.
Despite the relative abundance of case law IR datasets, case law retrieval remains an active and challenging area of research. Recent benchmarks report low baseline performance even when using forgiving metrics such as recall@1000 (CLERC) (Hou et al., 2024) or constructing queries only from text containing direct quotations from the relevant citations (LePARD) (Li and Harada, 2022).
2.4 Legal Contract Datasets
Several datasets have been published to facilitate AI research on legal contracts, but none address specific IR needs.
Prior datasets focus on clause extraction rather than retrieval. Chalkidis et al. (2017) & Leivaditi et al. (2020) provide datasets for extracting basic contract and lease information, not legal clauses, making them less useful for contract drafting. Hendrycks et al., 2021 introduces CUAD, a large-scale expert-annotated dataset with over 2,000 clauses across 41 categories from 510 commercial contracts. While useful for drafting, CUAD lacks ranking, an essential component of IR datasets. MAUD (Wang et al., 2023), ContractNLI (Koreeda and Manning, 2021), and LegalBench (Guha et al., 2023) focus on reading comprehension and reasoning, but would need extensive expert annotations for IR research. BigLaw Bench (Core, Workflow, and Retrieval) (Harvey AI, 2024) targets day-to-day legal tasks but is not specific to contract drafting and has an order of magnitude fewer annotations than ACORD.
ACORD seeks to provide a domain-specific expert-annotated large-scale IR benchmark for contract clause retrieval.
3 ACORD: A Contract Clause Retrieval Dataset
Clause Category Example Query
Limitation of Liability a party’s liability for fraud, negligence, personal injury or tort subject to a cap
Indemnification indemnification of first party claims
Affirmative Covenants insurance clause
Restrictive Covenants no solicit of customers
Term & Termination termination for convenience
Governing Law clause with multiple governing laws
Table 2: A subset of the 114 unique queries created by legal experts to address diverse contract drafting scenarios. Each query corresponds to a specific clause category, such as “Limitation of Liability.” The full list of the clause categories and queries is in Section C.3.
Clause Retrieval.
Clause retrieval involves identifying the most relevant precedent clauses within contracts based on a given query. This process ensures that lawyers can efficiently reference and adapt precedent clauses to meet specific drafting needs.
Task Definition.
ACORD frames the retrieval of clauses as an ad hoc information retrieval task. Given an attorney-written query and a corpus of clauses, search systems output a list of clauses by order of predicted relevance to the query. Performance is a function of the ordered lists (sometimes called a “run”) and the attorney’s ground-truth ratings. We measure performance with established metrics, such as the standard NDCG score and a task-specific normalized precision score.
Queries and Clauses.
The dataset includes 114 unique queries written by legal experts to address diverse drafting requirements. Each query targets one or more legal concepts across 9 clause categories. Sample queries are shown in Table 2, with a complete list provided in Section C.3.
Query-Clause Pair Score.
Each query-clause pair is assigned a 1-5 point relevance score. Clauses rated 3-5 stars are relevant, 2-star clauses are not relevant but helpful (from the same category, e.g., indemnification), and 1-star clauses are irrelevant. A sample-scored query-clause pair is shown in Table B-1, and the annotation rubrics are in Table A-1.
For each query, annotators aim to find 10 relevant clauses (3-5 stars) and 20 2-star clauses. Some queries have fewer than 10 relevant clauses, but all include at least 5. To avoid false negatives, 1-star clauses are selected from expert-annotated clauses in CUAD (Hendrycks et al., 2021), and 2-5 star Limitation of Liability and Indemnification clauses serve as 1-star for other categories. Though 1-star clauses are not individually rated, they are irrelevant due to being from distinct categories. Table C-2 shows the statistics of rated clauses, and Table C-3 lists the number of clauses by rating for each category.
Since most queries have fewer than five 5-star or 4-star clauses, x-star precision@5 scores are normalized to a 0-1 scale to adjust for the limited availability of high-rated clauses, as explained in Section F.1. When calculating the precision, the scores are rounded: 4.666 rounds up to 5, 4.333 rounds down to 4, etc.
Contract Corpus and Clause Corpus.
About 400 of such contracts are from CUAD  (Hendrycks et al., 2021), which was sourced from the Electronic Data Gathering, Analysis, and Retrieval (EDGAR) system maintained by the U.S. Securities and Exchange Commission (SEC). 50 contracts are Terms of Services, ToS, published online by selected Fortune 500 companies. We call these EDGAR contracts and online ToS the “Contract Corpus.” Annotators extracted from the Contract Corpus all clauses responsive to the 9 categories, which we call the “Clause Corpus.” Annotators then extracted clauses from the Clause Corpus responsive to each of the 114 queries. When extracting a clause responsive to a query, annotators include the entire subsection or section to ensure comprehensive context and understanding. As shown in Figure C-1, clauses in ACORD vary in length, with over half of the 3,000 clauses rated 2- through 5- stars having over 100 words. To assess the models’ retrieval performance, ACORD simplifies the real-life task of clause retrieval, which would require the models first to extract the Clause Corpus from the Contract Corpus. The benchmark included in this paper uses the Clause Corpus only.
Data Splits
We split the Clause Corpus into train, validation, and test sets. We form these splits at the query level, randomly allocating 45%, 5%, and 50% of the queries to the train, validation, and test splits, respectively, while ensuring that at least one query from each category is represented in the test set.
Data Statistics
ACORD contains 114 unique queries across 9 clause categories. It has over 126,000 query-clause pairs, each rated with a 1-5 score. See Section C.3 for more details of the queries and expert-rated clauses.
No False Negatives
Information retrieval datasets often contain false negatives due to an inability to annotate the entire dataset; however, every clause in ACORD is annotated to avoid false negatives.
Labeling Process
The labeling process follows five steps: (1) extraction, (2) retrieval, (3) scoring, (4) reconciliation, and (5) expansion.
First (Extraction), student annotators extract relevant clauses for the 9 categories from the Contract Corpus. Student annotators received 5-10 hours of training on the Limitation of Liability and Indemnification clauses. For the other 7 categories, we reused expert-annotated clauses from CUAD. Second (Retrieval), annotators search the Clause Corpus to find 10 relevant and 20 2-star irrelevant clauses per query. Some queries yielded only 5-8 relevant clauses. Third (Scoring), two expert attorneys and one student assign a 3- to 5-star relevance score to each query-clause pair based on a detailed four-page guideline. Irrelevant clauses mistakenly identified as relevant receive a 2-star rating from attorneys. Fourth (Reconciliation), a panel of 3-6 experienced attorneys reviews and reconciles scores for pairs with disagreements exceeding 2 points or disputes over relevance. The panel adjusts scores to stay within 2 points, and the final score is the average of the three annotators’ ratings. Annotator disagreement is tracked by the number of averaged scores revised during reconciliation. Lastly (Expansion), irrelevant 1-star clauses are added to each query as described in the “Query-Clause Pair Score” section above.
Data Format
The ACORD dataset is released in BEIR format, with a modification to the format of the qrels files to account for our explicit 1-star ratings. See Section C.5 for details.
4 Experiments
4.1 Setup
Metrics
We use the following five standard information retrieval metrics to measure performance: NDCG@5, NDCG@10, 5-star precision@5, 4-star precision@5, 3-star precision@5. NDCG measures the models’ ranking quality by how well they rank the most relevant clauses at the top of the search results, focusing on the top results. The x-star-precision@5 metrics evaluate the precision within the top 5 results by counting how many of the top 5 results meet or exceed the respective relevance score. See Section F.1 for details and definitions of the performance metrics.
Baseline Models
We evaluate several stand-alone retriever methods, including BM25 (Robertson and Zaragoza, 2009), MiniLM bi-encoder (22.7M parameters) (Wang et al., 2020) and OpenAI text embedding (OpenAI, 2024b). BM25 is a computationally inexpensive lexical or keyword-based retriever. In BM25, documents have higher scores if they contain terms from the query. Terms that are rarer in the overall corpus have higher weighting, so they contribute more to the document score. MiniLM bi-encoder (66M parameters) is an embedding model trained via knowledge distillation from BERT on various embedding datasets, including MSMARCO.
We evaluate two-phase retrieval-reranker systems by reranking the top $100$ BM25 outputs with MiniLM cross-encoder (22.7M parameters), GPT-4o-mini, GPT-4o, Llama-3.2-1B (1.23B parameters), and Llama-3.2-3B (3.21B parameters) (OpenAI, 2024a; Meta, 2024) models.
Fine-tuning
We finetune the MiniLM cross-encoder on the training dataset. We choose the learning rate and the number of updates via grid search on the validation NDCG@10 score. See Section F.2 for full training details.
4.2 Results
We show the results of most tested search systems on the test dataset in Table 3 with full results in Appendix D. Overall, a BM25 retriever paired with pointwise GPT-4o performs best, fine-tuned models perform better than models without fine-tuning, and embedding models perform better than lexical models. However, results vary significantly by query and model. Additionally, results using a pointwise reranking are generally much better than those using pairwise reranking in our experiments; see Table D-5 for a comparison.
NDCG@5 NDCG@10 3-star prec@5 4-star prec@5 5-star prec@5
Retriever Reranker
OpenAI embeddings (small) None 55.1 55.8 50.5 34.7 8.3
OpenAI embeddings (large) None 62.1 64.1 58.6 38.9 11.0
BM25 None 52.5 54.0 50.9 38.9 9.0
Cross-Encoder-MiniLM 59.3 60.9 60.0 43.5 6.2
GPT4o 76.9 79.7 81.1 60.0 17.2
GPT4o-mini 75.2 78.2 78.6 58.2 18.6
Llama-1B 13.8 14.4 13.0 10.5 4.1
Llama-3B 62.6 65.3 63.9 48.1 9.7
Table 3: Summary of the performance in % of selected retrieval models on the ACORD test dataset. The table reports NDCG@5, NDCG@10, and normalized precision@5 (for 3-star, 4-star, and 5-star clauses) aggregated across categories. We highlight the highest and second-highest metrics in bold and underlined, respectively. See Table D-3 in Section D.2 for results for all tested retriever methods. These results use a pointwise reranking as they are generally much better than the results using pairwise reranking in our experiments; See Table D-5.
Performance of Methods
In retrieval-only systems, the large OpenAI text embedding model performed best. Large LLM rerankers paired with the bi-encoder retriever significantly outperformed other methods. The GPT4o reranker with bi-encoder-miniLM achieved the highest results, with an NDCG@5 of 79.1% and an NDCG@10 of 81.2%. The GPT4o reranker with BM25 was a close second, outperforming all other methods by a noticeable margin. Smaller LLMs performed much worse than larger ones; for instance, Llama 3.2 showed a difference of 40%-points in most metrics between the 1B and 3B models.
Performance after Fine-Tuning
We finetuned the MiniLM cross-encoder on the training data and observed a modest increase of 2.0% and 5.4% in NDCG@5 and NDCG@10 scores, respectively. Meanwhile, the 3- and 4-star precisions decrease slightly, with a 5.1% increase in 5-star precision. Refer to Figure 2 for results before and after fine-tuning, and Table D-2 in Section D.1 for results by category. We also fine-tune the cross-encoder when paired with the bi-encoder for the initial retrievals; the results are in Table D-4.
Figure 2: The performance of BM25 with the MiniLM cross-encoder before and after fine-tuning. We see improvements in NDCG and 5-star precision scores. There are minor declines in the 3-star and 4-star precision scores. These results underscore the importance of domain-specific fine-tuning for advancing clause retrieval capabilities; however, the improvements are not large, underlining that the problem is not solved. See Table D-2 in Section D.1 for results for each category.
Performance by Clause Categories and Queries.
We measure the performance of the models across each of the 9 clause categories and under each query and find performance varies substantially. Some results are unexpected. For example, the query “indemnity of broad-based claims” achieved perfect scores for NDCG@5 and 3-star and 4-star precision@5 using BM25 paired with MiniLM cross-encoder, whereas “unilateral liability cap” and “as-is clause” got zero NDCG@5 score. This contradicts what we observed from human annotators: extracting “unilateral liability cap” and “as-is clauses” was easier than extracting relevant “indemnity of broad-based claims” clauses. Results for each clause category can be found in Section D.1.
5 Discussion and Future Work
Lexical Matching Difficulties
Some queries perform better with the MiniLM bi-encoder retriever than with BM25. For example, BM25 scores 0 on all metrics for the “as-is clause” query (see Table E-2) and near-zero scores even when paired with GPT-4o as the reranker. The reranking does not help much because BM25 does not retrieve relevant clauses in the top 100 results; the reranker only operates on the top 100 results. Hence, the recall@100 would be low or even zero. The dense retriever MiniLM bi-encoder performs much better than BM25 on the “as-is clause" query and many other queries.
Strong Extraction Performance
The models have demonstrated a strong understanding of legal terminology beyond simple lexical matching. For instance, in response to the query “change of control”, the models effectively return clauses referencing “ownership changes” or the “sale of substantially all assets”. Similarly, for the query “IP ownership assignment/transfer,” the models accurately identify “work for hire” clauses. The query “indemnification of broad-based claims” generates 100% NDCG@5 and 3-star and 4-star precision@5 scores, showcasing deep semantic comprehension (broad-based claims mean claims brought by a contracting party or any third party for breach of contract and tort, etc.). This contributes to the overall promising NDCG@5 and 3-star precision@5 scores, indicating the models’ relative maturity in clause extraction tasks.
Poor Performance on Legal Jargon without Context
Despite the strong capability to understand legal terminologies, the models perform poorly on queries of legal jargon without context. For example, the query “as-is clause” achieves zero NDCG@5 score across several models, whereas a simple keyword search would yield near-perfect results. The query “IP ownership assignment/transfer” returns joint ownership clauses instead of straight-forward ownership assignment clauses. In contrast, models perform significantly better with longer, context-rich queries. The query “IP infringement indemnity that covers trademark or copyright” achieves near-perfect scores across several models. To further test the effectiveness of more contextualized queries, we expand two queries (“as-is clause” and “change of control”) into medium and long formats. We find that medium formats yield significantly better results, as shown in Appendix E. In the Supplemental Materials, we include expanded medium and long formats for each of the 114 queries. We invite the research community to experiment further using these formats as prompts.
Low Ranking Performance
Models consistently fail to retrieve 4- or 5-star clauses in the top 5 results, with the 4-star and 5-star precision@5 scores being only 60.0% and 17.2% for the overall best method (BM25 with GPT-4o pointwise reranking). Many 4- or 5-star clauses are ranked beyond the top 10 or even 15. Low-ranking performance diminishes user confidence and limits the broader adoption of these models for real-world contract-related tasks. This finding is not surprising, given that there is currently no expert-annotated IR dataset for contracts, and the gap highlights the need for large-scale expert-annotated IR datasets like ACORD.
NDCG Is Not an Optimal Metric
The high NDCG scores seen do not adequately reflect the quality of the models in real-life contract drafting tasks. In the context of clause retrieval, the more meaningful metrics are 4- and 5-star precision@5 scores. This is because ranking relevant clauses requires a high degree of legal expertise. Most users of legal IR applications may lack such expertise. Thus, presenting many 3-star clauses as high-quality results can mislead users into relying on poorly drafted precedents, ultimately resulting in low-quality contracts.
The best method, bi-encoder retriever with GPT4o pointwise reranking, had an NDCG@5 score of 79.1%, but the 4-star and 5-star precision@5 scores were only 62.1% and 17.2%, respectively. Moreover, 40% of the queries tested have a 4-star precision@5 score below 50%. With ACORD, we hope the research community can advance in developing retrieval methodologies prioritizing clause quality.
Pointwise Outperforms Pairwise
We test ACORD using pointwise and pairwise approaches on different LLMs with the results available in Table D-5. Interestingly, the pointwise approach performs better than the pairwise approach for all tested retrieval methods, except when using the smallest Llama 3.2 model as the reranker. The same result persists after we deploy several methodologies in our pairwise experiments, including Pairwise Ranking Prompting used by Qin et al. (2024). This result seems inconsistent with prior literature, which generally indicates better results for pairwise methods (Qin et al., 2024; Zhuang et al., 2024). We encourage further research to validate and expand on our findings.
6 Conclusion
We introduced ACORD, the first information retrieval dataset for contract drafting fully annotated by experts. ACORD aims to facilitate research on clause retrieval, a foundational contract drafting task focusing on the most complex and heavily negotiated clauses in commercial contracts. ACORD includes 114 queries and over 126,000 query-clause pairs, each ranked on a scale from 1 to 5 stars. We evaluated 20 retrieval methods on ACORD and found that performance is promising but still has significant room for improvement. We also showed that model design markedly influences performance, suggesting that algorithmic improvements from the NLP community will help solve this challenge.
7 Limitations
ACORD is a simplified dataset because it contains a small scope of clauses already extracted from contracts. For example, ACORD does not have clauses about representatives and warranties, product descriptions, or payment terms. The performance will deteriorate in real-life legal tasks that would require the models to first extract relevant clauses from a large number of contracts before ranking them. We plan to publish the underlying contracts to enable research and experiments on the extraction and retrieval tasks.
Expert annotator ratings largely depend on a lawyer’s individual experience, industry, and interpretation of the query, among other things. This subjectivity is reflected in the annotators’ disagreement rate of 21%. However, although the ratings may not be exact, the relative ranking among different query-clause pairs is less uncertain.
The vast majority of commercial contracts are confidential and proprietary. ACORD uses publicly available sources, namely contracts in EDGAR and selected online terms of services. Although ACORD may not fully represent an enterprise’s database of negotiated contracts, the types of queries and clauses it includes are fairly standardized, making ACORD a valuable resource for enterprise contract retrieval.
ACORD focuses primarily on English-language contracts for U.S.-based companies, which limits its applicability to contracts governed by other legal systems or written in other languages.
Expert lawyers generate queries in ACORD and likely do not reflect the queries that a less experienced non-legal professional would have asked. Translating these queries into plain English would require legal expertise. This highlights the importance of keeping human legal professionals in the loop when creating products using ACORD.
Appendix A Appendix – Annotation Details
Rating Label Description
1 Not relevant, not useful The clause is not responsive to the query at all.
2 Not relevant, but useful The clause is not responsive to the query but contains language that could be helpful in drafting.
3 Relevant, but not perfect The clause is responsive to the query but has some defects. Such defects include, among other things, the clause: (1) is too long, (2) is missing key concepts typically seen in a real-life contract clause, (3) does not use standard legal language typically seen in a real-life contract, (4) is too one-sided and unfair, or (5) is unclear, etc.
4 Perfect The clause is relevant, concise, clear, and covers all necessary elements of the query.
5 Exemplary The clause meets all criteria for a 4 rating AND includes additional helpful legal concepts that may be responsive to the query.
Table A-1: ACORD Annotation rubric. Ratings range from 1 (not relevant) to 5 (exemplary), with each level defined by clear criteria.
A.1 Annotator Instructions and Interface
Annotator Instructions
Annotators were provided the five-point relevance rubric in table A-1 and the following instructions:
Please rate each clause assigned to you on a scale of 1-5 based on the following rubric. Clauses receiving a “Yes” for Relevance should receive a 3-5 rating. Clauses receiving a “No” for Relevance should receive a 1-2 rating.
Annotation Interface
Annotators complete Google Forms that include the query-clause pairs by rating each pair with a 3 to 5 star (or 2 if the clause is irrelevant). Annotators used Google Sheet to annotate the 2-star rated clauses. Results from completed Google Forms are then consolidated into the master dataset.
A.2 Annotator Recruitment and Demographics
The annotators were recruited by a nonprofit organization for the purpose of creating ACORD. The lawyer annotators are volunteers and do not receive compensation. Student annotators receive hourly compensation. All annotators were informed of the purpose of their involvement and consented to the inclusion of their annotations in ACORD.
Out of the annotators who consented to the disclosure of their demographic information, 42% identify as male and 58% identify as female. Except for one annotator based in the United Kingdom, all annotators are based in the United States.
Appendix B Appendix – Query-Clause Pair with Score
Sample of an annotated query-clause pair.
Query Clause Score
IP infringement exception to indirect damage waiver 8 INDEMNIFICATION.
8.1 By Commerce One. Commerce One shall indemnify, defend and hold harmless Corio and its Customers from any and all damages, liabilities, costs and expenses (including reasonable attorneys’ fees) incurred by Corio or its Customers arising out of any claim that the software infringes any patent, copyright, trademark or trade right secret of a third party; <omitted>
9 LIMITATION OF LIABILITY.
EXCEPT FOR LIABILITY ARISING UNDER SECTION 8 OF THIS AGREEMENT, IN NO EVENT SHALL EITHER PARTY’S LIABILITY ARISING OUT OF THIS AGREEMENT OR THE USE OR PERFORMANCE OF THE SOFTWARE EXCEED THE TOTAL AMOUNT ACTUALLY PAID BY CORIO HEREUNDER FOR THE TRANSACTION WHICH THE LIABILITY RELATES TO DURING THE TWELVE (12) MONTHS IMMEDIATELY PRIOR TO THE FILING OF THE CAUSE OF ACTION TO WHICH THE LIABILITY RELATES. EXCEPT FOR LIABILITY ARISING UNDER SECTION 8 OF THIS AGREEMENT, IN NO EVENT SHALL EITHER PARTY HAVE ANY LIABILITY TO THE OTHER PARTY FOR ANY LOST PROFITS OR COSTS OF PROCUREMENT OF SUBSTITUTE GOODS OR SERVICES, OR FOR ANY INDIRECT, SPECIAL OR CONSEQUENTIAL DAMAGES HOWEVER CAUSED AND UNDER ANY THEORY OF LIABILITY AND WHETHER OR NOT SUCH PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGE. THE PARTIES AGREE THAT THIS SECTION 9 REPRESENTS A REASONABLE ALLOCATION OF RISK.
4.67
Table B-1: A sample query-clause pair with a score averaged from three expert annotations.
Appendix C Appendix – Data Details
C.1 Licensing
ACORD is licensed under CC-BY-4.0.
C.2 Ethics Board Review
This research project was reviewed for ethics considerations by the the ETH Zürich Ethics Commission.
C.3 Data List and Statistics
This subsection presents some details of the data. LABEL:tab:_categories_and_queries shows all 114 queries and the categories they belong to. While Tables C-2 and C-3 show summary statistics of the number of clauses for each query split by rating and counts for the number of clauses for each category and rating.
Table C-1: The table lists the 9 clause categories and the 114 queries in ACORD.
No. Clause Category Query
1 Limitation of Liability cap on liability
2 Limitation of Liability liability cap is based on purchase price
3 Limitation of Liability precedents for insurance coverage influencing limitation of liability
4 Limitation of Liability Fix fee liability cap
5 Limitation of Liability Cap on liability equals 12 months payment
6 Limitation of Liability cap on liability for indirect damages
7 Limitation of Liability unilateral liability cap
8 Limitation of Liability mutual liability cap
9 Limitation of Liability two parties having different liability caps and/or carveouts
10 Limitation of Liability Cap on liability without carveouts
11 Limitation of Liability liability cap carveouts
12 Limitation of Liability cap on liability subject to law
13 Limitation of Liability compliance with law carveout to cap on liability
14 Limitation of Liability indemnification carveout to cap on liability
15 Limitation of Liability IP infringement exception to cap on liability
16 Limitation of Liability third party IP infringement exception to cap on liability
17 Limitation of Liability liability cap clauses that exclude third party IP infringement and fraud, gross negligence or willful misconduct
18 Limitation of Liability personal or bodily injury exception to liability cap
19 Limitation of Liability confidentiality exceptions to liability cap
20 Limitation of Liability fraud, negligence or willful misconduct carveout to liability cap
21 Limitation of Liability a party’s liability for fraud, negligence, personal injury or tort subject to a cap
22 Limitation of Liability personal or bodily injury exception to cap on liability via indemnification carveout
23 Limitation of Liability third party IP infringement exception to cap on liability via indemnification carveout
24 Limitation of Liability fraud, gross negligence or willful misconduct exception to cap on liability via indemnification carveout
25 Limitation of Liability seller-favorable cap on liability clauses
26 Limitation of Liability buyer-favorable cap on liability clauses
27 Limitation of Liability non-reliance clause
28 Limitation of Liability as-is clause
29 Limitation of Liability as-is clause with carveouts
30 Limitation of Liability unqualified “as-is” clause
31 Limitation of Liability warranty disclaimer clause that includes implied warranties
32 Limitation of Liability warranty disclaimer clause that disclaims implied warranties of merchantability and fitness for a particular purpose
33 Limitation of Liability waiver of implied warranty of title and non-infringement
34 Limitation of Liability warranty disclaimer that does not specifically waive title and non-infringement warranties
35 Limitation of Liability exclusive remedy for breach of product warranty
36 Limitation of Liability product replacement, repair or refund as exclusivity remedy
37 Limitation of Liability product warranty of shelf life
38 Limitation of Liability product warranty around manufacturing and shipping
39 Limitation of Liability customer’s right for defective products
40 Limitation of Liability seller-favorable warranty disclaimer clauses
41 Limitation of Liability buyer-favorable warranty disclaimer clauses
42 Limitation of Liability disclaimer of indirect damages
43 Limitation of Liability consequential damages waiver
44 Limitation of Liability incidental damages disclaimer
45 Limitation of Liability disclaimer of lost profits
46 Limitation of Liability disclaimer of punitive damages
47 Limitation of Liability disclaimer of strict liability
48 Limitation of Liability unilateral indirect damages waiver
49 Limitation of Liability mutual indirect damages waiver
50 Limitation of Liability indirect damage waiver is subject to law
51 Limitation of Liability indemnification carveout to indirect damage waiver
52 Limitation of Liability IP infringement exception to indirect damage waiver
53 Limitation of Liability third party IP infringement exception to indirect damage waiver
54 Limitation of Liability indirect damage waiver clauses that exclude third party IP infringement and fraud, gross negligence or willful misconduct
55 Limitation of Liability personal or bodily injury exception to indirect damage waiver
56 Limitation of Liability confidentiality exceptions to indirect damage waiver
57 Limitation of Liability fraud, gross negligence or willful misconduct carveout to indirect damage waiver
58 Limitation of Liability indirect damages waiver applies to a party’s liability for fraud, negligence or personal injury
59 Limitation of Liability personal or bodily injury exception to waiver of indirect damages via indemnification carveout
60 Limitation of Liability first party claim exception to waiver of indirect damages
61 Limitation of Liability fraud, gross negligence or willful misconduct exception to waiver of indirect damages via indemnification carveout
62 Limitation of Liability seller-favorable waiver of indirect damages clauses
63 Limitation of Liability buyer-favorable waiver of indirect damages clauses
64 Indemnification indemnity or indemnification clause
65 Indemnification mutual indemnification provisions
66 Indemnification unilateral indemnification clause
67 Indemnification indemnification of third party claims
68 Indemnification fraud and/or gross negligence indemnity
69 Indemnification third party IP infringement indemnity
70 Indemnification indemnification of third-party claims based on breach of agreement
71 Indemnification IP infringement indemnity that covers trademark or copyright
72 Indemnification indemnification covers indirect claims
73 Indemnification Indemnification of first party claims
74 Indemnification indemnification clauses that include hold harmless
75 Indemnification indemnification clauses that do not include hold harmless
76 Indemnification indemnification clauses that include defending claims
77 Indemnification indemnification clause that allows indemnifying party to control defenses
78 Indemnification Indemnity of broad-based claims
79 Indemnification Indemnified party includes affiliates
80 Indemnification first party indemnification of specified claims
81 Indemnification first party indemnification of broad-based claims
82 Indemnification indemnification of broad-based third party claims
83 Indemnification third party claim indemnity of fraud, negligence or willful misconduct
84 Indemnification Third Party claim indemnity limited to use of products in compliance with agreement
85 Indemnification indemnification that covers violation of law
86 Indemnification seller-favorable indemnification clauses
87 Indemnification buyer-favorable indemnification clauses
88 Affirmative Covenants Revenue/Profit Sharing
89 Affirmative Covenants Minimum Commitment
90 Affirmative Covenants Audit Rights
91 Affirmative Covenants Insurance
92 Restrictive Covenants Most Favored Nation Clause
93 Restrictive Covenants Non-compete bound by time and territory
94 Restrictive Covenants Exclusivity bound by time and territory
95 Restrictive Covenants No-Solicit of Customers
96 Restrictive Covenants No-Solicit Of Employees not bound by time or longer than 12 months
97 Restrictive Covenants Non-Disparagement
98 Restrictive Covenants Rofr/Rofo/Rofn
99 Restrictive Covenants Change Of Control
100 Restrictive Covenants Anti-Assignment clause that requires notice only for assignment to affiliates
101 Restrictive Covenants Covenant Not To Sue
102 Term Renewal clause that requires notice to Renew
103 Term Clause that requires notice to terminate auto-renew
104 Term Termination for Convenience
105 Governing Law New York Governing Law
106 Governing Law England Governing Law
107 Governing Law Clause with multiple governing laws
108 Governing Law Governing Law excluding UCC or other similar regulatory frameworks
109 Liquidated Damages Liquidated Damages
110 Third party beneficiary Third Party Beneficiary
111 IP ownership/license IP Ownership Assignment or Transfer
112 IP ownership/license Joint IP Ownership
113 IP ownership/license License clause covering affiliates of licensor and/or licensee
114 IP ownership/license Source Code Escrow
Score Min Mean Median Std. Dev. Max
1-star 395 1080.42 1293.00 383.28 1293
2-star 15 20.13 20.00 2.50 28
3-star 1 4.11 4.00 2.45 17
4-star 1 5.82 6.00 2.48 12
5-star 1 2.17 2.00 1.45 6
Table C-2: The min, mean, median, standard deviation, and max statistics of the number of rated clauses for each score. Refer to Table C-3 for the number of query-clause pairs in the data for each of the 9 clause categories.
Number of Queries Number of ratings in the dataset Total ratings
Category 5-star 4-star 3-star 2-star 1-star
Limitation of Liability 63 87 345 200 1277 81459 83368
Indemnification 24 29 153 95 459 31032 31768
Affirmative Covenants 4 4 27 17 80 1581 1709
Restrictive Covenants 10 3 64 45 206 3956 4274
Term 3 0 20 12 66 1187 1285
Governing Law 4 2 19 10 82 1582 1695
Liquidated Damages 1 0 6 4 21 395 426
Third-party beneficiary 1 0 8 3 20 395 426
IP ownership/license 4 5 21 17 84 1581 1708
Total 114 130 663 403 2295 123168 126659
Table C-3: The table lists the number of queries and annotated clauses under each x-star precision@5 rating for each clause category in ACORD.
C.4 Clause Extraction Methodology in Clause Corpus
Most clauses are extracted at a section or subsection level, marked by a numerical header such as “Section 9” or “8(a)”. If a single sentence within a section or subsection is relevant to the query, the entire section or subsection is included. Additionally, if the section or subsection references another section, such other section is also included. An <omitted> symbol is inserted between them to indicate that the two sections are not continuous in the contract. Clauses have widely varying lengths, ranging from 13 words to 1,898 words. Figure C-1 shows the distribution of 2- through 5-star clause lengths in ACORD.
Figure C-1: Histogram over the number of words in 2- to 5-star clause in the ACORD dataset. The figure also includes statistics showing that 2- to 5-star clauses have an average of 225 words and a median of 146 words in each clause.
C.5 Data Format
We release the ACORD dataset in BEIR format, with a modification to the format of the qrels files to account for explicit annotation of 1-star judgements.
C.5.1 BEIR format
In the BEIR format for ad-hoc IR datasets, the clauses and queries can be found in corpus.jsonl and queries.jsonl, respectively. The query-clause scores for the train, development, and test splits can be found in qrels/train, test, valid.tsv. As mentioned in the dataset section, lawyer judgments are one-indexed (between one and five stars), but it’s desirable to have zero-indexed scores when calculating NDCG. Therefore, in the qrels files, we subtract one from each lawyer’s judgment to get scores that range from zero to four.
C.5.2 Explicit zero-scores
We deviate from the standard BEIR format in that we explicitly, rather than implicitly, encode irrelevant query-clause pairs. In the standard BEIR format, irrelevant query-clause pairs (with a score of zero) are omitted from the qrels files, which implies their irrelevance. As with all Crenshaw- or TREC-style IR datasets where unjudged query-clause pairs are treated as irrelevant, this can lead to many false negatives (Thakur et al., 2021). In ACORD qrel files, we explicitly label irrelevant clauses with a score of zero. Omitted query-clause pairs should be ignored during evaluation because their score is unknown.
In constructing ACORD, we carefully constructed our dataset to reduce the incidence of false negatives. We explicitly label irrelevant clauses through lawyer judgments and through selecting clause sources that are known to be irrelevant to particular queries.
Appendix D Appendix – Model Performance
The next two sections will show detailed and extended results for the model evaluation of the test data.
D.1 Model Performance by Category
NDCG@5 NDCG@10 3-star precision@5 4-star precision@5 5-star precision@5
Category
Limitation of Liability 0.750 0.785 0.771 0.557 0.200
Indemnification 0.763 0.803 0.800 0.600 0.178
Affirmative Covenants 0.930 0.910 1.000 0.867 0.000
Restrictive Covenants 0.682 0.662 0.850 0.650 0.200
Term 0.912 0.936 1.000 0.600 NaN
Governing Law 0.805 0.868 0.800 0.500 0.200
Liquidated Damages 0.812 0.788 0.800 0.600 NaN
Third-party beneficiary 1.000 0.974 1.000 1.000 NaN
IP ownership/license 0.764 0.787 0.867 0.600 0.067
Table D-1: The table shows the performance by clause category when using BM25 as the retriever model paired with pointwise GPT4o for reranking. We get Not a Number, NaN, for the 5-star precison@5 if there are no 5-star clauses for the queries in the category. Cells show the mean over test queries in a given category where we ignore NaN values.
In this section, we break the model performance down by category to determine how the different categories influence the performance of the models. We show in Table D-1 the results for BM25 with GPT4o as the reranker model. Overall, we see that the 5-star precision@5 is struggling across all categories, but in particular, with “Affirmative Covenants” and “IP ownership/license”, resulting in a low overall mean. However, for “Affirmative Covenants,” the NDCG scores are some of the highest across categories, thus showing that NDCG might not be the best metric in this setting as discussed in Section 5. Table D-2 shows the results by category before and after fine-tuning the cross-encoder on the training data. Overall, the improvement by fine-tuning the data is limited, which could indicate that more data or better architectures are needed.
NDCG@5 NDCG@10 3-star precision@5 4-star precision@5 5-star precision@5
Category Fine-tuned
Limitation of Liability No 0.541 0.565 0.507 0.364 0.043
Indemnification 0.721 0.727 0.743 0.529 0.089
Affirmative Covenants 0.605 0.571 0.667 0.533 0.000
Restrictive Covenants 0.566 0.539 0.750 0.500 0.000
Term 0.558 0.684 0.400 0.200 NaN
Governing Law 0.517 0.544 0.600 0.500 0.200
Liquidated Damages 0.838 0.765 1.000 0.600 NaN
Third-party beneficiary 0.887 0.941 1.000 0.800 NaN
IP ownership/license 0.382 0.458 0.333 0.333 0.067
Limitation of Liability Yes 0.642 0.694 0.598 0.448 0.138
Indemnification 0.678 0.735 0.638 0.500 0.081
Affirmative Covenants 0.373 0.403 0.444 0.222 0.133
Restrictive Covenants 0.565 0.509 0.717 0.567 0.067
Term 0.447 0.583 0.200 0.133 NaN
Governing Law 0.655 0.768 0.633 0.367 0.200
Liquidated Damages 0.245 0.328 0.333 0.133 NaN
Third-party beneficiary 0.841 0.859 0.933 0.533 NaN
IP ownership/license 0.424 0.499 0.422 0.222 0.067
Table D-2: The performance by clause category of the BM25 retriever model with the cross-encoder reranker model both with and without fine-tuning the cross-encoder on the training data. The finetuning hyperparameters were tuned by maximizing the mean NDCG@5 score for the validation set. We get Not a Number, NaN, for the 5-star precison@5 if there are no 5-star clauses for the queries in the category. Cells show the mean over all test queries in a given category where we ignore NaN values.
D.2 Model Performance and Size for All Tested Models
In this subsection, we show the results for all tested models using the test data. The results can be seen in Table D-3 and show that the best and second best rerankers are GPT4o and GPT4o-mini when using either BM25 or Bi-Encoder-MiniLM. For the Llama-based rerankers, we see that the 3B size model far outperforms the 1B model. Thus, the model size has a large impact both when using BM25 and when using Bi-Encoder-MiniLM as the initial retriever.
NDCG@5 NDCG@10 3-star prec@5 4-star prec@5 5-star prec@5 Size
Retriever Reranker
OpenAI embeddings (small) None 0.551 0.558 0.505 0.347 0.083 N/A
OpenAI embeddings (large) None 0.621 0.641 0.586 0.389 0.110 N/A
BM25 None 0.525 0.540 0.509 0.389 0.090 0M
Cross-Encoder-MiniLM 0.593 0.609 0.600 0.435 0.062 66M
GPT4o 0.769 0.797 0.811 0.600 0.172 1.8T
GPT4o-mini 0.752 0.782 0.786 0.582 0.186 8B
Llama-1B 0.138 0.144 0.130 0.105 0.041 1.23B
Llama-3B 0.626 0.653 0.639 0.481 0.097 3.21B
Bi-Encoder-MiniLM None 0.571 0.572 0.498 0.358 0.076 66M
Cross-Encoder-MiniLM 0.601 0.610 0.586 0.428 0.069 132M
GPT4o 0.791 0.812 0.814 0.621 0.172 1.8T
GPT4o-mini 0.763 0.790 0.775 0.586 0.200 8B
Llama-1B 0.200 0.205 0.147 0.119 0.048 1.29B
Llama-3B 0.628 0.657 0.628 0.481 0.110 3.27B
Table D-3: The table shows the results for all tested models on the test data. For each of the retrieval models BM25 and Bi-Encoder-MiniLM, we bold and underline the best and second best scores, respectively. We also include the size of the models. For GPT4o and GPT4o-mini, the exact sizes are unknown when writing, but we use estimates found online Howarth (2024). For the OpenAI embedding models, the exact sizes are unknown, and we did not find estimates.
D.3 Model Performance after Fine-tuning
We show below in Table D-4 the results of fine-tuning the cross-encoder reranker to the training data before evaluating the methods on the test data. We tested with both BM25 and bi+encoder as the initial retriever models. Overall, fine-tuning improves the models slightly; however, the results are still worse than using GPT4o or GPT4o-mini as the rerankers, cf. Table D-3.
NDCG@5 NDCG@10 3-star prec@5 4-star prec@5 5-star prec@5
Retriever Reranker
Not fine-tuned
BM25 Cross-Encoder-MiniLM 59.3 60.9 60.0 43.5 6.2
Bi-Encoder-MiniLM Cross-Encoder-MiniLM 60.1 61.0 58.6 42.8 6.9
Fine-tuned
BM25 Cross-Encoder-MiniLM 61.3 (+2.0) 66.3 (+5.4) 59.4 (-0.6) 43.3 (-0.2) 11.3 (+5.1)
Bi-Encoder-MiniLM Cross-Encoder-MiniLM 64.7 (+4.6) 68.1 (+7.1) 61.6 (+3.0) 44.9 (+2.1) 12.2 (+5.3)
Table D-4: Performance in % before and after fine-tuning the cross-encoder with pointwise reranking. We write the change in each metric after fine-tuning in parenthesis. We test both with BM25 and bi-encoder as the retriever model. Overall, the results show a modest improvement in NDCG scores and 5-star precision, as well as 3- and 4-star precisions for the bi-encoder model. However, with BM25, we see a slight decline in 3- and 4-star precisions.
D.4 Model Performance for Pairwise and Pointwise Rerankings
We show below in Table D-5 the results when using pairwise reranking rather than pointwise reranking. We tested it with GPT4o, GPT4o-mini, Llama 3.2 1B, and Llama 3.2 3B. Overall, we see that the results are much worse when using pairwise reranking. However, Llama 3.2 1B does see a significant improvement in its metrics, but it is still worse than the other models when they use pointwise reranking.
NDCG@5 NDCG@10 3-star prec@5 4-star prec@5 5-star prec@5
Retriever Reranker
Pointwise
BM25 GPT4o 76.9 79.7 81.1 60.0 17.2
GPT4o-mini 75.2 78.2 78.6 58.2 18.6
Llama-1B 13.8 14.4 13.0 10.5 4.1
Llama-3B 62.6 65.3 63.9 48.1 9.7
Pairwise
BM25 GPT4o 58.0 59.1 56.1 43.9 13.1
GPT4o-mini 57.4 58.5 55.4 43.5 12.4
Llama-1B 52.1 54.3 50.2 37.9 9.0
Llama-3B 52.8 54.4 50.5 38.9 9.0
Table D-5: Performance in % of GPT4o, GPT4o-mini, Llama 3.2 1B, and Llama 3.2 3B reranker when using either pointwise or pairwise reranking with BM25 as the retriever model. Overall, the pairwise results are much worse than those for pointwise reranking. However, Llama 3.2 1B does see a significant benefit from the more computationally heavy technique.
Appendix E Appendix – Modifying Query Format
In this section, we present results for modifying the query to use a medium and long format rather than the default short queries. An expert lawyer constructs the medium and long-format queries. We show in Table E-1 the results using different formats, demonstrating that providing more context to the models significantly improve performance. Although we only tested the “Change of Control” and “as-is” queries, in the Supplemental Materials, we provide medium and long format queries for all 114 queries and encourage the research community to conduct further experiments to improve the models’ performance.
Query Type Variation Query
Change of control Short “Change Of Control”
Medium “Clause that prohibits change of control itself”
Long “Does one party have the right to terminate or is consent or notice required of the counterparty if such party undergoes a change of control, such as a merger, stock sale, transfer of all or substantially all of its assets or business, or assignment by operation of law?”
’as-is’ Short “’as-is’ clause”
Medium “’as-is’ clause” that disclaims all warranties’
Long “Is this a clause that states that the goods or services are being provided on an ’as-is’ basis in their current condition, with no warranties or guarantees regarding their quality, performance, or suitability?”
Table E-1: Query variations for two query types: “Change of control” and “as-is”. Each variation represents a different length or level of detail, with queries ranging from short, concise statements to long, detailed descriptions.
NDCG@5 NDCG@10 3-star precision@5 4-star precision@5 5-star precision@5
Model Query Type Variation
BM25 Change of control Short 0.914 0.856 0.800 0.600 0.200
Medium 0.887 0.794 1.000 0.600 0.200
Long 0.470 0.470 0.200 0.200 0.200
’as-is’ Short 0.000 0.000 0.000 0.000 0.000
Medium 0.683 0.660 0.400 0.400 0.200
Long 0.284 0.297 0.200 0.200 0.000
BM25 + GPT-4o Change of control Short 0.775 0.821 1.000 0.400 0.200
Medium 0.449 0.561 0.600 0.000 0.000
Long 0.835 0.850 1.000 0.200 0.200
’as-is’ Short 0.102 0.068 0.000 0.000 0.000
Medium 1.000 0.954 1.000 1.000 0.000
Long 1.000 0.917 1.000 0.800 0.000
BM25 + Cross-Encoder Change of control Short 0.832 0.770 0.800 0.600 0.200
Medium 0.590 0.707 0.400 0.200 0.000
Long 0.629 0.717 0.400 0.400 0.000
’as-is’ Short 0.000 0.000 0.000 0.000 0.000
Medium 0.887 0.794 1.000 0.600 0.200
Long 0.518 0.641 0.600 0.200 0.000
Table E-2: Comparison of retrieval performance across different query variations, categories, and models. The table reports five quality metrics (NDCG@5, NDCG@10, 3-star precision@5, 4-star precision@5, and 5-star precision@5) for each model (BM25, BM25 + GPT-4o, and BM25 + Cross-Encoder) evaluated on three different formats for the “Change of Control” and “as-is” queries. Using the more detailed queries, the medium or long formats, shown in Table E-1, give better results. However, the long format is not always an improvement over the medium format.
Appendix F Appendix – Experiment and Evaluation Details
F.1 Performance Metrics
Performance is measured using standard metrics like NDCG@5, NDCG@10, 5-star-precision@5, 4-star-precision@5, and 3-star-precision@5. The NDCG@k measures the normalized discounted cumulative gain of the top k returned results. We use the standard definition of NDCG@k seen in (1).
$DCG@k$ $=\sum_{i=1}^{k}\frac{rel_{i}}{\log_{2}(i+1)},$
$IDCG@k$ $=\sum_{i=1}^{|REL_{k}|}\frac{rel_{i}}{\log_{2}(i+1)},$
$NDCG@k$ $=\frac{DCG@k}{IDCG@k},$ (1)
where $rel_{i}$ is the graded relevance of the document at position $i$, and $REL_{k}$ is the list of relevant documents sorted by relevance. We use the pytrec_eval library to calculate the NDCG scores.
The k-star precision@5 precision metric counts the number of returned clauses in the top 5 results rated $\geq k$. If there are less than 5 viable clauses for a query, i.e., clauses with ratings $\geq k$, then we normalize by the number of viable clauses. This ensures the precision metrics always range from 0 to 1. Formally, the k-star precision@5 is defined in (2).
[…]
where $\mathbbm{1}$ is the indicator function so $\mathbbm{1}(rel_{i}\geq k)$ is $1$ if $rel_{i}\geq k$ and 0 otherwise.
F.2 Fine-tuning Experiments
We perform a grid search over the learning rate in $\{5\times 10^{-5},1\times 10^{-4},3\times 10^{-4},5\times 10^{-4},1\times 10^{-3}\}$ and number of updates in $\{100,200,400,600,800,1000,1500,2000,2500,3000\}$, choosing the hyperparameters with the best average validation NDCG@10 score over three runs.
We oversampled the data so that an equal number of examples of each relevance appear in the training data. We use a batch size of $64$ and the AdamW optimizer with a weight decay of $0.01$. We use cross-entropy loss, where the probability targets are the normalized relevance scores. For example, a one-star clause corresponds to a target of $p=0.0$, and a four-star clause corresponds to a target of $p=0.75$.
F.3 Compute
Fine-tuning experiments and Llama evaluations were performed on an NVIDIA RTX A6000 GPU. For fine-tuning experiments the full grid-search took about three GPU hours. Evaluation of bi-encoder and cross-encoder MiniLM models took less than five minutes per run. Llama pointwise evaluations took about thirty minutes per model, and pairwise evaluations took about an hour per model.